OpenStudy (anonymous):

the three sides of an equilateral triangle measure 12x - 12, 6x + 12, and 3x + 24. What is the legnth of each side of the triangle? Please explain how to solve!!!

OpenStudy (anonymous):

what do we know about equilateral triangle ?

OpenStudy (anonymous):

all sides have same length so we can set them equal to each other 12x - 12= 6x +12 6x= +24 x=4 plug into one sides 12x - 12 48-12=36 *notes we should get same number if we plug into different expression

OpenStudy (anonymous):

Imran just did you homework sonja :P

OpenStudy (anonymous):

@FoolforMath thats all it gave me! :P Thankyou both though i have exams tommorow and this is my study guide and i just couldnt figure that problem out for the life of me. xD

OpenStudy (anonymous):

only for this problem, we don't give out answer; at least I don't